OpenStudy (anonymous):

Need help on graphing inequalities 18x-9y<2

OpenStudy (anonymous):

just take that 18x-9y=2... draw the line, check either sides of the the graph wherever you got this function true will be your solution set... samjhi?

OpenStudy (anonymous):

But what are the points in the graph where they supposed to be

OpenStudy (anonymous):

OKAY , divide the whole equation by 2 , so that RHS has 1 , u get 18x/2 - 9y/2 = 1

OpenStudy (anonymous):

its upto your choice.. we just have to check it on either side

OpenStudy (anonymous):

9x - 9/2y = 1 plug in values of x and find the value of y , plot the graph

OpenStudy (anonymous):

oh , wait 9x - 9/2y < 1 now whatever is x & y's coefficient , make it 1 , by dividing for x , it is 9 , for y it is 9/2 , put them on x & y axis respectively

OpenStudy (anonymous):

draw a dotted line through the points , shade the region away from zero as the inequality is greater than 1